Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by razaul karim » Tue May 14, 2013 12:42 pm
From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1

(B) 0.2

(C) 0.25

(D) 0.4

(E) 0.6
Ans:
This means that P(M and N both selected) = [spoiler](2/5) x (1/4) = 1/10[/spoiler]

Another question::
Joshua and Jose work at an auto repair center with 4 other workers. For a survey on health care insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

a)1/15
b)1/12
c)1/9
d)1/6
e)1/3

Ans:
[spoiler](1/6)*(1/5)+(1/6)*(1/5)=2/30=1/15[/spoiler]
Here why do we consider Both the options (Jose*Joshuna) and (Joshuna*jose)
But in the first question we considerded probability just 2/5x1/4 and why not (2/5x1/4)(1/4x2/5) like we did in the second problem.Please explain the difference..
Please explain
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Tue May 14, 2013 12:50 pm
razaul karim wrote:From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1

(B) 0.2

(C) 0.25

(D) 0.4

(E) 0.6
Ans:
This means that P(M and N both selected) = [spoiler](2/5) x (1/4) = 1/10[/spoiler]

Another question::
Joshua and Jose work at an auto repair center with 4 other workers. For a survey on health care insurance, 2 of the 6 workers will be randomly chosen to be interviewed. What is the probability that Joshua and Jose will both be chosen?

a)1/15
b)1/12
c)1/9
d)1/6
e)1/3

Ans:
[spoiler](1/6)*(1/5)+(1/6)*(1/5)=2/30=1/15[/spoiler]
Here why do we consider Both the options (Jose*Joshuna) and (Joshuna*jose)
But in the first question we considerded probability just 2/5x1/4 and why not (2/5x1/4)(1/4x2/5) like we did in the second problem.Please explain the difference..
Please explain
The second question can be solved using the exact same approach used in the first question.
That is, P(Joshua and Jose both chosen) = (2/6) x (1/5) = 1/15

The solution you posted still has the correct answer (1/15). The only difference is that it considers two acceptable outcomes:
1) Joshua chosen 1st and Jose chosen 2nd
2) Jose chosen 1st and Joshua chosen 2nd

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by sanaa.rizwan » Thu May 23, 2013 6:29 am
I am not getting the solution. why is the probability of the first occurrence/selection 2/5 and not 1/5

When 2 events occur one after the other, we are suppose to multiply their probabilities.

The probability that Marnie is selected is 1/5
The probability that Noomi is selected is 1/4 (since only 4 employees are left)

so the probability that both Marnie and Noomi are selected is 1/5 * 1/4 = 1/20

This is the method used in one of the MGMAT practise test explanations.

Where am I going wrong?
Join the discussion

by Brent@GMATPrepNow » Thu May 23, 2013 6:38 am
sanaa.rizwan wrote:I am not getting the solution. why is the probability of the first occurrence/selection 2/5 and not 1/5

When 2 events occur one after the other, we are suppose to multiply their probabilities.

The probability that Marnie is selected is 1/5
The probability that Noomi is selected is 1/4 (since only 4 employees are left)

so the probability that both Marnie and Noomi are selected is 1/5 * 1/4 = 1/20

This is the method used in one of the MGMAT practise test explanations.

Where am I going wrong?
Our goal is to find P(M and N both selected)

There are two ways to approach this.

Method #1:
P(M and N both selected) = P(one of them is selected 1st AND the other selected 2nd)
= P(one of them is selected 1st) x P(the other selected 2nd)
= (2/5)(1/4)
= 1/10
Aside: P(one of them is selected 1st) = 2/5 because I'm allowing for either Marnie or Noomi to be selected first.

Method #2:
P(M and N both selected) = P(M selected 1st AND N selected 2nd OR N selected 1st AND M selected 2nd)
= P(M selected 1st AND N selected 2nd) + P(N selected 1st AND M selected 2nd)
= (1/5)(1/4) + (1/5)(1/4)
= 1/20 + 1/20
= 1/10

Both solutions yield the same answer.

Your approach, seems to consider one possible outcome (M selected 1st AND N selected 2nd) and doesn't consider the other one (N selected 1st AND M selected 2nd)

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by sanaa.rizwan » Thu May 23, 2013 6:43 am
Now I get it why is is 2/5 and not 1/5.

Thanks a lot
Join the discussion

by Scott@TargetTestPrep » Thu Jan 04, 2018 8:40 am
razaul karim wrote:From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1

(B) 0.2

(C) 0.25

(D) 0.4

(E) 0.6
There are a total of 5C2 = 5!/(3! x 2!) = (5 x 4)/2 = 10 ways of choosing two people from a group of five people. Since the selection of Marnie and Noomi corresponds to one of these choices, the probability of this selection is 1/10 = 0.1.

Answer: A

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion